OpenStudy (danjs):

maximize the volume, from a given surface area... @katie00

OpenStudy (anonymous):

??

OpenStudy (danjs):

that second practice prob you liked me

OpenStudy (danjs):

What is the maximum volume for a sealed cylinder, if the surface area is set at 600pi

OpenStudy (danjs):

for a cylinder , volume and surface area \[V = \pi*r^2*h\] \[S.A = 2*\pi * r^2 + 2*\pi*r*h\]

OpenStudy (danjs):

given S.A is 600pi, what is the max possible value for the Volume...

OpenStudy (danjs):

One thing to think of, the graph of the function for A ' , the first derivative of Volume will show you when it is increasing/decreasing, depending on the sign of the derivative

OpenStudy (danjs):

Volume and S.A. are both functions of height, and radius, cant do the derivative of A with that

OpenStudy (danjs):

\[S.A = 600\pi = 2*\pi * r^2 + 2*\pi*r*h \] you can solve that surface area either for r or h, solving for h will be easier, since not roots to deal with h = (300/r) - r

OpenStudy (danjs):

put that h in terms of r, intto the Volume Function, substitute \[V = \pi*r^2*[\frac{ 300 }{ r } - r]\]

OpenStudy (danjs):

now you can differentiate that and get the rate of change of the Volume with respect to the radius \[V(r) = 300*\pi*r - \pi*r^3\] and \[V '(r) = 300\pi - 3*\pi*r^2\]

OpenStudy (danjs):

V ' will = 0 at r = 10

OpenStudy (danjs):

before r=10, 0<r<10 the Area is increasing after r=10, starts decreasing local max there

OpenStudy (danjs):

Overall - You want maximum Volume, that keeps in line with the surface area staying at 600pi 'optimization probs'

OpenStudy (danjs):

1 - put the volume in terms of one variable, we did r 2 - differentiate 3 - the maximums change from increasing functions to decreasing, that occurs at r = 10

OpenStudy (danjs):

most probs are like that process, just different things, boxes, and whjatever

OpenStudy (anonymous):

Ty I think I understadn this one much better than the previous one you did. I was working on the previous one and it's still kind of confusing.

OpenStudy (danjs):

oh, can you link me back there

OpenStudy (danjs):

or we can start again new on it

OpenStudy (anonymous):

let's start a new one on that one plz!

OpenStudy (anonymous):

@DanJS would there be a grpah to this problem?